What is the appropriate z a value for a 75.4 % confidence interval for u? Round your response to at least 2 decimal places. N2
Answer #1

solution

At 75.4% confidence level

\alpha = 1 - 0.754

\alpha = 1 - 0.754 =0.246

\alpha/2 = 0.123

Z\alpha/2 = Z0.123 = + / -1.16